OpenStudy (anonymous):

Solve for x: |2x + 6| − 4 = 20 x = 9 and x = 11 x = −9 and x = 15 x = 9 and x = −15 No solution

OpenStudy (nurali):

|2x + 6| − 4 = 20 |2x + 6| =20+4 |2x + 6| =24 2x+6=24 or 2x+6=-24 2x=24-6 or 2x=-24-6 2x=18 or 2x=-30 x=9 or x=-15